OpenStudy (anonymous):

Which equations have a negative solution? Choose all answers that are correct. A. B. C. D.

OpenStudy (anonymous):

\[A.1/4w=-2\]

OpenStudy (anonymous):

B.\[-3=x/[-7]\]

OpenStudy (anonymous):

C.\[6=-y/5\]

OpenStudy (anonymous):

D.\[-2/3z=-8\]

OpenStudy (lolacole12):

b or A, But I am not sure. I am not good in this subject at all. I hate dealing with negatives :\

OpenStudy (anonymous):

its a multiple answer question

OpenStudy (lolacole12):

b or a

OpenStudy (lolacole12):

@Deadly_Roses

OpenStudy (lolacole12):

@megax or @rhettwalker14

OpenStudy (lolacole12):

@rhettwalker14 can u help

OpenStudy (anonymous):

no i dont understand this sorry

OpenStudy (anonymous):

oh wait it was A and C

OpenStudy (deadly_roses):

nvm looks like you got it already lol

OpenStudy (lolacole12):

oh i said a and b its a and c i was close